OpenStudy (anonymous):

What is the solution to the equation 4^( –2x) ≈ 3 ? x = –0.631 x = –0.396 x = 0.396 x = 0.631

OpenStudy (anonymous):

\[x =\frac{\ln 3}{-2\ln{4}}\] Entering this on your calculator, you should get x = -0.396 So Answer is B